# Simple Sublinear Algorithms for $(\Delta+1)$ Vertex Coloring

### v3

@@ 5 Concluding Remarks
In this paper, we simplified the sublinear algorithms of [ACK19] for $(\Delta+1)$ vertex coloring of graphs with maximum degree $\Delta$. This was achieved by weakening the palette sparsification theorem (PST) of [ACK19] to our asymmetric palette sparsification theorem (APST), which we show admits a much simpler proof and coloring algorithm, without limiting its applicability to sublinear algorithms. At this point, there are several natural questions for future work, which we elaborate on below.
@@
Since its introduction in [ACK19], PST has found various applications in designing sublinear algorithms for other graph coloring problems as well. It will be interesting to examine to what extent APST is tailored to $(\Delta+1)$ coloring and whether, similar to PST, APST can also be extended to other graph coloring problems. For instance, [AA20] developed sublinear algorithms for $O(\Delta/\log(\Delta))$ coloring of triangle-free graphs using a different PST targeted to this problem. Can we also design an APST for this problem?
@@
Another set of questions are regarding “white-box” applications of PST. For instance, [AKM22] showed how to use ideas from PST, plus various other tools, to design a graph streaming algorithm for $\Delta$-coloring. Can we use APST to obtain a similar result but without the complicated algorithms and elaborate case analyses in [AKM22]? Taking this even further, can one obtain sublinear algorithms for $(\Delta-1)$ coloring [Ree99] or even less colors as in [MR14] wherein the barrier to colorability are “small” subgraphs [^11] These results are known in the graph theory literature as strengthening of Brook’s theorem for $\Delta$-coloring which identifies having a $(\Delta+1)$-clique as the only barrier to $\Delta$ colorability for $\Delta\geqslant 3$. A highly general form of these results is that of [MR14] that shows the only barriers to $\Delta-k+1$ coloring, for large enough $\Delta$, are $O(\Delta)$-size subgraphs that are not $(\Delta-k+1)$ colorable, as long as $k$ satisfies $(k+1)\cdot(k+2)\leqslant\Delta$, and that this threshold is tight [EWHK98].. Another example of these white-box applications is in designing Local Computation Algorithms (LCA) for $(\Delta+1)$ coloring [CFG+19]; again, can APST lead to simpler and more efficient LCAs for $(\Delta+1)$ coloring, e.g., one with only $O(\Delta\cdot\log{(n)})$ query time (existing algorithms in [CFG+19] have some unspecified $\mbox{\rm poly}(\Delta)$ dependence)? Yet another example, is designing dynamic graph algorithms for $(\Delta+1)$ coloring, against adaptive adversaries [BRW25, FH25] that use sparse-dense decompositions and somewhat similar coloring steps as in the PST. Again, can ideas from APST simplify and further improve these bounds? It is worth mentioning that already some tools developed in this paper are used in [BRW25] for addressing this problem.
@@
Finally, beside its applicability to sublinear algorithms, PST has also been studied extensively from a purely combinatorial point of view [AA20, ABD22, KK23, KK24, HK24, AK24]. It thus is natural to consider APST through the same lens as well; for instance, in the same vein as in [HK24], we can ask the following question: suppose we sample lists of average size only two (or some other larger constant) colors on vertices chosen from some range $\left\{1,\ldots,q\right\}$; then, what is a (asymptotically) minimal choice for $q$ so that the greedy algorithm that colors the vertices in the increasing order of their list sizes finds a proper coloring of any given graph from its sampled lists with high probability? For instance, does $q=\Theta(\Delta^{2})$ work in this case (perhaps for $\Delta$ sufficiently large or even $\omega(\log{n})$)? An alternative version of this question, which to our knowledge is not studied even for PST, is to fix a choice of $q$ to be, say, $\Delta+1$, and instead determines what fraction of vertices can be properly colored this way?
@@ Acknowledgement
We would like to thank Soheil Behnezhad, Yannic Maus, and Ronitt Rubinfeld for helpful discussions and their encouragement toward finding a simpler sublinear algorithms for $(\Delta+1)$ vertex coloring. We are also thankful to Chris Trevisan for pointing out a shorter proof of Lemma A.5 included in this version of the paper. Sepehr Assadi is additionally grateful to Yu Chen and Sanjeev Khanna for their prior collaboration in [ACK19] that formed the backbone of this work. Finally, we are grateful to the anonymous reviewers of TheoretiCS for many invaluable comments and suggestions that helped tremendously with the presentation of the paper.
Part of this work was conducted while the first named author was visiting the Simons Institute for the Theory of Computing as part of the Sublinear Algorithms program.
@@ Appendix A Appendix: A Self-contained Sublinear Time Algorithm
In addition to our main result, we also present a very simple sublinear time algorithm for $(\Delta+1)$ coloring in the following theorem. This algorithm has recently and independently been discovered in [FHC25] and has also appeared as part of lecture notes for different courses [Lec24, Lec25, HW24] at this point. Finally, this algorithm can also be seen as a simple adjustment to the sublinear time $(\Delta+o(\Delta))$ coloring algorithm of [MS21].
Theorem 4.
@@
We note that unlike our sublinear time algorithm in Theorem 3 which was non-adaptive, namely made all its queries in advance before seeing the answer to them, the current algorithm is adaptive and needs to receive the answer to each query before deciding its next query.
The algorithm in Theorem 4 is quite similar to the standard greedy algorithm for $(\Delta+1)$ coloring. It iterates over the vertices and colors each one greedily by finding a color not used among the neighbors of this vertex yet (which exists by pigeonhole principle). However, unlike the greedy algorithm, $(i)$ it crucially needs to iterate over the vertices in a random order, and, $(ii)$ instead of iterating over the neighbors of the current vertex to find an available color, it samples a color randomly for this vertex and then iterates over all vertices with this color to make sure they are not neighbor to the current vertex. Formally, the algorithm is as follows.
Algorithm 1.
An (adaptive) sublinear time algorithm for $(\Delta+1)$ vertex coloring. 1. Let $C_{1},C_{2},\ldots,C_{\Delta+1}$ be the color classes to be output at the end, initially set to empty. 2. Pick a permutation $\pi$ of vertices in $V$ uniformly at random. 3. For $v\in V$ in the order of the permutation $\pi$: (a) Sample $c\in[\Delta+1]$ uniformly at random. (b) For every vertex $u\in C_{c}$, check if $(u,v)$ is an edge in $G$; if Yes, restart from Line (3a). (c) If the algorithm reaches this step, color the vertex $v$ with $c$ and add $v$ to $C_{c}$.
@@
For any input graph $G=(V,E)$, the expected runtime of 1 is
$O(\frac{n^{2}}{\Delta}\cdot\log{\Delta})~\text{time}.$
To continue, we need to set up some notation. For any vertex $v\in V$, we define the random variable $X_{v}$ as the number of $(u,v)$ queries checked by the algorithm in the for-loop of coloring $v$. Additionally, for any $v\in V$ and permutation $\pi$ picked over $V$, define $N^{<}_{\pi}(v)$ as the neighbors $u$ of $v$ with $\pi(u)<\pi(v)$, namely the ones that are colored before $v$ by Algorithm 1. Let $\deg^{<}_{\pi}(v):=\left|{N^{<}_{\pi}(v)}\right|$. We start with the basic observation that the runtime of 1 can be stated in terms of the variables $\left\{X_{v}\right\}_{v\in V}$.
Observation A.2.
The expected runtime of 1 is $O(\sum_{v\in V}\operatorname*{{\mathbb{E}}}\left[X_{v}\right])$.
Proof.
@@
Our task is now to bound each of $\operatorname*{{\mathbb{E}}}\left[X_{v}\right]$ for $v\in V$ to bound the runtime of the algorithm using A.2.
@@
Claim A.4.
Fix any vertex $v\in V$, any choice of the permutation $\pi$, and any assignment of colors $C(u_{1}),C(u_{2}),\ldots$ by 1 to all vertices that appear before $v$ in $\pi$. Then,
@@
[…]
Using Claim A.4, we can conclude the proof of Lemma A.3.
Proof of Lemma A.3.
Each color $c^{\prime}\in[\Delta+1]$ is chosen with probability $1/(\Delta+1)$ in Line (3a) of the algorithm. Thus,
@@
Using Claim A.4 and the above bounds, we conclude
@@
[…]
By Lemma A.3 (and A.2), for any choice of the permutation $\pi$ in 1,
@@
$\operatorname*{{\mathbb{E}}}_{\pi}\left[\sum_{v\in V}\frac{n}{\Delta+1-\deg^{<}_{\pi}(v)}\right]=O(\frac{n^{2}}{\Delta}\cdot\log{\Delta}).$
Proof.
@@
$\operatorname*{{\mathbb{E}}}_{\pi}\left[\frac{1}{\Delta+1-\deg^{<}_{\pi}(v)}\right]$ $=\sum_{d=0}^{\deg(v)}\frac{1}{\deg(v)+1}\cdot\frac{1}{\Delta+1-d}\leqslant\sum_{d=0}^{\Delta}\frac{1}{\Delta+1}\cdot\frac{1}{\Delta+1-d}=O(\frac{\log{\Delta}}{\Delta}),$
@@
Lemma A.1 now follows immediately from Eq 1 and Lemma A.5. We can now use this to wrap up the proof of Theorem 4.
Proof of Theorem 4.
First, we consider the case $\Delta\geqslant\sqrt{n\log n}$. In this case, we use 1, which relies on the adjacency matrix to access the input graph. By Lemma A.1 we know that this algorithm has expected runtime $O(\frac{n^{2}}{\Delta}\cdot\log{\Delta})$ which is $O(n\sqrt{n\cdot\log(n)})$ by the lower bound on $\Delta$.

### v4

@@ 5 Concluding Remarks
In this paper, we simplified the sublinear algorithms of [undeff] for $(\Delta+1)$ vertex coloring of graphs with maximum degree $\Delta$. This was achieved by weakening the palette sparsification theorem (PST) of [undeff] to our asymmetric palette sparsification theorem (APST), which we show admits a much simpler proof and coloring algorithm, without limiting its applicability to sublinear algorithms. At this point, there are several natural questions for future work, which we elaborate on below.
@@
Since its introduction in [undeff], PST has found various applications in designing sublinear algorithms for other graph coloring problems as well. It will be interesting to examine to what extent APST is tailored to $(\Delta+1)$ coloring and whether, similar to PST, APST can also be extended to other graph coloring problems. For instance, [undefa] developed sublinear algorithms for $O(\Delta/\log(\Delta))$ coloring of triangle-free graphs using a different PST targeted to this problem. Can we also design an APST for this problem?
@@
Another set of questions are regarding “white-box” applications of PST. For instance, [undefg] showed how to use ideas from PST, plus various other tools, to design a graph streaming algorithm for $\Delta$-coloring. Can we use APST to obtain a similar result but without the complicated algorithms and elaborate case analyzes in [undefg]? Taking this even further, can one obtain sublinear algorithms for $(\Delta-1)$ coloring [undefag] or even less colors as in [undefad] wherein the barrier to colorability are “small” subgraphs [^8] These results are known in the graph theory literature as strengthening of Brook’s theorem for $\Delta$-coloring which identifies having a $(\Delta+1)$-clique as the only barrier to $\Delta$ colorability for $\Delta\geqslant 3$. A highly general form of these results is that of [undefad] that shows the only barriers to $\Delta-k+1$ coloring, for large enough $\Delta$, are $O(\Delta)$-size subgraphs that are not $(\Delta-k+1)$ colorable, as long as $k$ satisfies $(k+1)\cdot(k+2)\leqslant\Delta$, and that this threshold is tight [undefo].. Another example of these white-box applications is in designing Local Computation Algorithms (LCA) for $(\Delta+1)$ coloring [undefl]; again, can APST lead to simpler and more efficient LCAs for $(\Delta+1)$ coloring, e.g., one with only $O(\Delta\cdot\log{(n)})$ query time (existing algorithms in [undefl] have some unspecified $\mbox{\rm poly}(\Delta)$ dependence)? Yet another example, is designing dynamic graph algorithms for $(\Delta+1)$ coloring, against adaptive adversaries [undefi, undefu] that use sparse-dense decompositions and somewhat similar coloring steps as in the PST. Again, can ideas from APST simplify and further improve these bounds? It is worth mentioning that already some tools developed in this paper are used in [undefi] for addressing this problem.
@@
Finally, beside its applicability to sublinear algorithms, PST has also been studied extensively from a purely combinatorial point of view [undefa, undefb, undefaa, undefab, undefy, undefc]. It thus is natural to consider APST through the same lens as well; for instance, in the same vein as in [undefy], we can ask the following question: suppose we sample lists of average size only two (or some other larger constant) colors on vertices chosen from some range $\left\{1,\ldots,q\right\}$; then, what is a (asymptotically) minimal choice for $q$ so that the greedy algorithm that colors the vertices in the increasing order of their list sizes finds a proper coloring of any given graph from its sampled lists with high probability? For instance, does $q=\Theta(\Delta^{2})$ work in this case (perhaps for $\Delta$ sufficiently large or even $\omega(\log{n})$)? An alternative version of this question, which to our knowledge is not studied even for PST, is to fix a choice of $q$ to be, say, $\Delta+1$, and instead determines what fraction of vertices can be properly colored this way?
@@ Acknowledgement
We would like to thank Soheil Behnezhad, Yannic Maus, and Ronitt Rubinfeld for helpful discussions and their encouragement toward finding a simpler sublinear algorithms for $(\Delta+1)$ vertex coloring. We are also thankful to Chris Trevisan for pointing out a shorter proof of Lemma A.7 included in this version of the paper. Sepehr Assadi is additionally grateful to Yu Chen and Sanjeev Khanna for their prior collaboration in [undeff] that formed the backbone of this work. Finally, we are grateful to the anonymous reviewers of TheoretiCS for many invaluable comments and suggestions that helped tremendously with the presentation of the paper.
@@ Appendix A Appendix: A Self-contained Sublinear Time Algorithm
In addition to our main result, we also present a very simple sublinear time algorithm for $(\Delta+1)$ coloring in the following theorem. This algorithm has recently and independently been discovered in [undefq] and has also appeared as part of lecture notes for different courses [undefd, undefah] at this point. Finally, this algorithm can also be seen as a simple adjustment to the sublinear time $(\Delta+o(\Delta))$ coloring algorithm of [undefae].
Theorem A.1.
@@
We note that unlike our sublinear time algorithm in Theorem 4.1 which was non-adaptive, namely made all its queries in advance before seeing the answer to them, the current algorithm is adaptive and needs to receive the answer to each query before deciding its next query.
The algorithm in Theorem A.1 is quite similar to the standard greedy algorithm for $(\Delta+1)$ coloring. It iterates over the vertices and colors each one greedily by finding a color not used among the neighbors of this vertex yet (which exists by pigeonhole principle). However, unlike the greedy algorithm, $(i)$ it crucially needs to iterate over the vertices in a random order, and, $(ii)$ instead of iterating over the neighbors of the current vertex to find an available color, it samples a color randomly for this vertex and then iterates over all vertices with this color to make sure they are not neighbor to the current vertex. Formally, the algorithm is as follows.
1
2Let $C_{1},C_{2},\ldots,C_{\Delta+1}$ be the color classes to be output at the end, initially set to empty. ;
3
4Pick a permutation $\pi$ of vertices in $V$ uniformly at random.;
5
6For $v\in V$ in the order of the permutation $\pi$:
7
8  Sample $c\in[\Delta+1]$ uniformly at random. ;
9   For every vertex $u\in C_{c}$, check if $(u,v)$ is an edge in $G$; if Yes, restart from Line (1). ;
10   If the algorithm reaches this step, color the vertex $v$ with $c$ and add $v$ to $C_{c}$. ;
11
12
Algorithm 1 An (adaptive) sublinear time algorithm for $(\Delta+1)$ vertex coloring.
@@
For any input graph $G=(V,E)$, the expected runtime of Algorithm 1 is
$O\left(\frac{n^{2}}{\Delta}\cdot\log{\Delta}\right)~\text{time}.$
To continue, we need to set up some notation. For any vertex $v\in V$, we define the random variable $X_{v}$ as the number of $(u,v)$ queries checked by the algorithm in the for-loop of coloring $v$. Additionally, for any $v\in V$ and permutation $\pi$ picked over $V$, define $N^{<}_{\pi}(v)$ as the neighbors $u$ of $v$ with $\pi(u)<\pi(v)$, namely the ones that are colored before $v$ by Algorithm Algorithm 1. Let $\deg^{<}_{\pi}(v):=\left|{N^{<}_{\pi}(v)}\right|$. We start with the basic observation that the runtime of Algorithm 1 can be stated in terms of the variables $\left\{X_{v}\right\}_{v\in V}$.
{observation}
The expected runtime of Algorithm 1 is $O(\sum_{v\in V}\operatorname*{{\mathbb{E}}}\left[X_{v}\right])$.
Proof A.3.
@@
Our task is now to bound each of $\operatorname*{{\mathbb{E}}}\left[X_{v}\right]$ for $v\in V$ to bound the runtime of the algorithm using Appendix A.
@@
Claim 3.
Fix any vertex $v\in V$, any choice of the permutation $\pi$, and any assignment of colors $C(u_{1}),C(u_{2}),\ldots$ by Algorithm 1 to all vertices that appear before $v$ in $\pi$. Then,
@@
$\operatorname*{{\mathbb{E}}}\left[X_{v}\mid\pi\right]=\frac{\operatorname*{{\mathbb{E}}}\left[\left|{C_{c}}\right|\mid\pi\right]}{\operatorname*{\textnormal{Pr}}\left(\text{$c$ does not appear in $N^{<}_{\pi}(v)$}\mid\pi,C(u_{1}),C(u_{2}),\ldots\right)}.$
Using Claim 3, we can conclude the proof of Lemma A.4.
Proof A.6 (Proof of Lemma A.4).
Each color $c^{\prime}\in[\Delta+1]$ is chosen with probability $1/(\Delta+1)$ in Line (1) of the algorithm. Thus,
@@
Using Claim 3 and the above bounds, we conclude
@@
$\leqslant\frac{n}{\Delta+1-\deg^{<}_{\pi}(v)}.$
By Lemma A.4 (and Appendix A), for any choice of the permutation $\pi$ in Algorithm 1,
@@
$\operatorname*{{\mathbb{E}}}_{\pi}\left[\sum_{v\in V}\frac{n}{\Delta+1-\deg^{<}_{\pi}(v)}\right]=O\left(\frac{n^{2}}{\Delta}\cdot\log{\Delta}\right).$
Proof A.8.
@@
$\operatorname*{{\mathbb{E}}}_{\pi}\left[\frac{1}{\Delta+1-\deg^{<}_{\pi}(v)}\right]$ $=\sum_{d=0}^{\deg(v)}\frac{1}{\deg(v)+1}\cdot\frac{1}{\Delta+1-d}\leqslant\sum_{d=0}^{\Delta}\frac{1}{\Delta+1}\cdot\frac{1}{\Delta+1-d}=O\left(\frac{\log{\Delta}}{\Delta}\right),$
@@
Lemma A.2 now follows immediately from Equation 1 and Lemma A.7. We can now use this to wrap up the proof of Theorem A.1.
Proof A.9 (Proof of Theorem A.1).
First, we consider the case $\Delta\geqslant\sqrt{n\log n}$. In this case, we use Algorithm 1, which relies on the adjacency matrix to access the input graph. By Lemma A.2 we know that this algorithm has expected runtime $O(\frac{n^{2}}{\Delta}\cdot\log{\Delta})$ which is $O(n\sqrt{n\cdot\log(n)})$ by the lower bound on $\Delta$.